OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1).

OpenStudy (anonymous):

Ok for the first one I will solve it for y to easily find the slope of that. y=-3x+y From this equation the slope is -3 since it is before the x. If the line asked for is perpendicular to it, it will have the opposite slope. Can you find the slope of the line from that?

OpenStudy (anonymous):

Honestly have no clue. @Venny

OpenStudy (anonymous):

The opposite of a negative is a positive, so what is the opposite of -3?

OpenStudy (anonymous):

3. @Venny

OpenStudy (anonymous):

Yes so you have y=3x+b so far you are also given the point (6,-1) so you plug this into the equation to find b(y intercept) -1=3(6)+b -1=18+b -19=b do you understand that?

OpenStudy (anonymous):

yes, i do. @Venny

OpenStudy (anonymous):

That means your equation will be y=3x-19 :)

OpenStudy (anonymous):

thank you!! :) @Venny